### v1

# Doc2SAR: A Synergistic Framework for High-Fidelity Extraction of Structure-Activity Relationships from Scientific Documents

Abstract
Extracting molecular structure-activity relationships (SARs) from scientific literature and patents is essential for drug discovery and materials research. However, this task remains challenging due to heterogeneous document formats and limitations of existing methods. Specifically, rule-based approaches relying on rigid templates fail to generalize across diverse document layouts, while general-purpose multimodal large language models (MLLMs) lack sufficient accuracy and reliability for specialized tasks, such as layout detection and optical chemical structure recognition (OCSR). To address these challenges, we introduce DocSAR-200, a rigorously annotated benchmark of 200 scientific documents designed specifically for evaluating SAR extraction methods. Additionally, we propose Doc2SAR, a novel synergistic framework that integrates domain-specific tools with MLLMs enhanced via supervised fine-tuning (SFT). Extensive experiments demonstrate that Doc2SAR achieves state-of-the-art performance across various document types, significantly outperforming leading end-to-end baselines. Specifically, Doc2SAR attains an overall Table Recall of 80.78% on DocSAR-200, exceeding end2end GPT-4o by 51.48%. Furthermore, Doc2SAR demonstrates practical usability through efficient inference and is accompanied by a web app.
1 Introduction
The extraction of molecular structure-activity relationships (SARs) from scientific literature and patents is fundamental for accelerating the discovery of novel drugs and advanced materials [1, 2, 3]. Accurate and efficient extraction of molecular activity data not only supports computational modeling but also significantly reduces innovation cycles in pharmaceutical and materials science research [4, 5, 6, 7]. Despite its critical importance, automated SAR extraction remains challenging due to the complexity and heterogeneity of scientific documents.
Two major issues currently impede progress in automated SAR extraction. Firstly, there is a notable lack of standardized benchmark datasets specifically designed to evaluate and advance methods for this complex, multimodal information extraction task. Secondly, existing automated extraction methods exhibit critical deficiencies: traditional rule-based methods employing predefined templates are inflexible and fail to generalize across diverse document formats [8, 9, 10]; while general-purpose multimodal large language models (MLLMs), though promising [11, 12], still fall short in specialized subtasks [13, 14, 15, 16]. As our failure analysis in Figure 1 demonstrates, the practical utility of monolithic, end-to-end GPT-4o [17] is severely limited by a cascade of compounding errors. This inherent brittleness highlights the need for a more sophisticated, modular approach.
To address these challenges, we make two primary contributions. First, we introduce DocSAR-200, a meticulously curated benchmark of 200 diverse scientific documents designed to provide a standardized testbed for evaluating SAR extraction methods. Second, to tackle the complexities presented by this benchmark, we propose Doc2SAR, a novel synergistic framework. This framework integrates domain-specific tools like optical chemical structure recognition (OCSR) and fine-tuned MLLMs to precisely solve complex sub-tasks, with robust rule-based methods for post-processing.
Our key contributions can be summarized as follows:
  1. 1.
    We introduce DocSAR-200, a comprehensive and rigorously annotated benchmark specifically tailored for evaluating and driving progress in automated SAR extraction.
  2. 2.
    We propose Doc2SAR, a novel synergistic framework that intelligently combines domain-specific tools and MLLMs enhanced via supervised fine-tuning (SFT) to precisely solve complex sub-tasks, demonstrating a superior approach to naive end-to-end models.
  3. 3.
    Extensive experiments validate the effectiveness of Doc2SAR, which achieves state-of-the-art performance by attaining an overall Table Recall of 80.78% on the DocSAR-200 benchmark, representing an improvement of 51.48% over the strongest end-to-end baseline.
  4. 4.
    Doc2SAR demonstrates strong practical applicability, efficiently processing over 100 PDFs per hour on a single RTX 4090 GPU. Additionally, it features an interactive web interface for easy visualization, verification, and manual refinement, detailed in Appendix E.
Figure 1: Motivation for a modular framework: An end-to-end GPT-4o model shows compounding errors across crucial stages of SAR extraction.
2 DocSAR-200 Benchmark
Figure 2: A representative ground truth annotation from the DocSAR-200 benchmark. For each document, molecular structures and their corresponding entries in activity tables are identified and linked, often across multiple pages, to create a structured record as ground truth in CSV format.
To systematically evaluate and advance automated extraction of SARs from scientific literature, we introduce DocSAR-200, a high-quality benchmark comprising 200 rigorously annotated scientific documents, including both patents and research articles. The benchmark includes complex, real-world cases that require linking molecules and tables across multiple pages, as illustrated by the representative example in Figure 2.
Figure 3: Distribution of SMILES and PDFs across sources. Top row: statistics at the molecules level; bottom row: statistics at the documents level. From left to right: comparison between patent and literature sources; breakdown across literature journals; and distribution across patent authorities.
The final DocSAR-200 dataset includes 3329 valid SMILES, of which 2664 are derived from patent documents and 665 from scientific literature. While patents contribute the majority of molecules, literature-sourced SMILES—primarily from journals such as ACS Med. Chem. Lett. and J. Med. Chem.—tend to exhibit higher semantic clarity and labeling consistency. Overall, the dataset spans 98 patents and 102 research articles, covering documents in English, Chinese, and Japanese, and encompassing a wide range of chemical domains and formatting styles.
Figure 3 provides a structured overview of dataset composition across two axes of analysis. Vertically, the top row aggregates statistics at the molecular (i.e., SMILES) level, while the bottom row reflects the distribution of source documents (i.e., PDFs). Horizontally, each column represents a different source-based perspective: the leftmost column contrasts patent and literature contributions, the middle focuses on literature journal breakdowns, and the rightmost depicts distributions across major patent offices. These statistics provide insight into the heterogeneity and coverage of the benchmark corpus, supporting downstream evaluation under diverse scenarios.
Detailed annotation methodology and quality control procedures are provided in Appendix B.
3 Doc2SAR Framework
Given an input PDF document $D$, our goal is to extract a set of relationships =$\mathcal{R}=\{(s_{k},a_{k})\}_{k=1}^{N}$, where $s_{k}$ is a molecule’s SMILES string and $a_{k}$ is its associated activity.
To achieve this, we propose Doc2SAR, a novel synergistic framework. Recognizing the inherent brittleness of monolithic, end-to-end MLLMs on this complex task, Doc2SAR employs a modular design that decomposes the problem into independently optimizable components. This architecture achieves superior interpretability and robustness through a synergistic integration of two key elements: specialized, high-precision domain-specific tools for deterministic tasks like OCSR, and powerful, MLLMs for nuanced contextual understanding, such as table extraction and coreference recognition.
As illustrated in Figure 4, the framework operates through three main stages: Layout Detection, Parallel Extraction, and Post-Processing, which are detailed in the following subsections.
Figure 4: The Doc2SAR framework. The PDF is parsed into layout segments; molecular images and tables are processed in parallel. The framework’s strength lies in its synergistic design, using specialized tools like OCSR and fine-tuned MLLMs for perception, and rule-based methods for logical integration.
3.1 Layout Detection
Given a raw PDF document $\mathcal{D}$, a YOLO-based [18] layout detection module first segments the document into a set of image blocks, =$\mathcal{B}=\{\mathcal{B}_{\text{mol}},\mathcal{B}_{\text{table}}\}$. Here, $\mathcal{B}_{\text{mol}}$ represents detected molecular images and $\mathcal{B}_{\text{table}}$ represents table regions. This detection function can be formally defined as:
$\mathcal{B}=f_{\text{det}}(\mathcal{D}).$
3.2 Parallel Extraction
The regions identified by the layout detector are then processed through several concurrent, specialized modules. These modules form the core neural perception engine of our framework.
OCSR.
Each molecular image block $I^{\text{mol}}_{k}\in\mathcal{B}_{\text{mol}}$ is processed by our proprietary OCSR tool. This neural model [19, 20] employs a Swin Transformer [21] as the image encoder $f_{\text{enc}}$ and a BART-style [22] autoregressive decoder $f_{\text{dec}}$ to generate the corresponding SMILES [23] string $s_{k}$. The process is defined as:
$s_{k}=f_{\text{ocsr}}(I^{\text{mol}}_{k})=f_{\text{dec}}(f_{\text{enc}}(I^{\text{mol}}_{k})).$
Molecular Coreference Recognition.
An MLLM agent extracts the textual identifier for each molecule. For each molecular image $I^{\text{mol}}_{k}$, the agent analyzes the surrounding layout context $C^{\text{ctx}}_{k}$ to identify the corresponding coreference ID $c^{\text{mol}}_{k}$. This operation is defined as:
$c^{\text{mol}}_{k}=f_{\text{coref}}(I^{\text{mol}}_{k},C^{\text{ctx}}_{k}).$
Table HTML Extraction.
Each table image $I^{\text{table}}_{j}$ is processed by an MLLM agent guided by a conditional prompt. This prompt instructs the model to first screen the table for relevant bioactivity keywords (e.g., $\text{IC}_{50}$). Only if such keywords are found does the model proceed to transform the table image into a structured HTML representation $H_{j}$. This integrated approach simplifies the architecture while leveraging the MLLM’s ability to perform both classification and extraction.
3.3 Post-processing
As the final stage of the framework, this module synthesizes the outputs from the parallel extraction streams. It performs a crucial cross-modal alignment using the molecular coreference ID as the pivotal key. Specifically, given a molecular pair $(s_{k},c^{\text{mol}}_{k})$ and a tabular entry $(c^{(j)}_{i},a^{(j)}_{i})$, the module employs fuzzy string matching to align their respective IDs. A match is considered valid only if its similarity score exceeds a predefined threshold $\delta$. Each successfully linked pair is then consolidated into a structured SAR record $(s_{k},a^{(j)}_{i})$ for downstream analysis.
3.4 SFT for Key Components
While our framework leverages established tools for Layout Detection and OCSR, the more nuanced tasks of Table HTML Extraction and Molecular Coreference Recognition rely on MLLMs. To overcome the limitations of general-purpose models on these tasks, we employ a SFT strategy to enhance their performance. This involves fine-tuning a single Qwen2.5-VL-7B-Instruct model [24] on a consolidated dataset that combines data for these two sub-tasks. The consolidated training data comprises:
  • Table Extraction Data: 10,000 examples randomly sampled from the PubTabNet [25] training dataset for converting table images to HTML.
  • Molecular Coreference Recognition Data: Our custom dataset of 6,615 instances, including 1,000 synthetic hard cases [26], for identifying molecular identifiers. The detailed construction methodology is provided in Appendix C.
The fine-tuning process was conducted on a multi-GPU server for 3 epochs, using a batch size of 1 and a learning rate of $2\times 10^{-6}$. We adopted rigorous validation protocols throughout the SFT process to ensure the resulting model’s robustness and reliability across both tasks.
4 Experiments
In this section, we present a series of experiments designed to systematically evaluate our proposed Doc2SAR framework. Our evaluation aims to answer two key research questions: (1) How effective is our synergistic framework approach compared to standard end-to-end methods for SAR extraction? (2) To what extent does SFT contribute to the performance gains?
4.1 Experimental Setup
Evaluation Metric.
The primary evaluation is conducted on our DocSAR-200 benchmark, as detailed in Section 2. We measure performance using Table Recall, a stringent metric that calculates the percentage of data rows where the molecule (identified by its correct SMILES) and all of its associated activity values are extracted and linked perfectly. This metric is chosen for its ability to holistically reflect the practical effectiveness of a method in delivering complete and accurate structured output.
Baseline Models.
We compare our approach against a suite of powerful, state-of-the-art MLLMs. The baselines include OpenAI’s GPT-4o [17], Google’s Gemini-2.5-flash [27], Meta’s Llama3.2-VL-90B-Instruct [28], and Alibaba’s Qwen2.5-VL [24] series, specifically the 72B-Instruct and 7B-Instruct versions. All models are evaluated in a zero-shot setting using carefully crafted prompts.
Implementation Details.
For our experiments, we accessed proprietary models like GPT-4o and Gemini-2.5-flash via their official APIs, while open-source models, including the Llama and Qwen series, were hosted locally from public repositories. To ensure a fair and robust comparison, the task-specific prompts used for all models were developed and optimized on a small, disjoint set of validation documents not present in any of the test sets. To promote reproducibility, the DocSAR-200 benchmark and all evaluation scripts will be made publicly available at https://github.com/Jesse-zjx/Doc2SAR.
4.2 Main Results: End-to-End vs. Doc2SAR Framework
This section presents the comprehensive evaluation of our framework on the DocSAR-200 benchmark, with detailed results shown in Table 1. The analysis is structured to first cover the overall performance, followed by granular breakdowns by document type and sub-task.
Table 1: Table Recall (%) on the DocSAR-200 benchmark.
Model Table Recall (%)
Overall Doc. Type Sub. Task
Patents Liters. Coref. Rec. Table. Ex.
# End to End Method
GPT-4o 29.30 20.75 39.19 0.00 49.23
Gemini-2.5-flash 23.03 12.32 40.27 0.00 38.54
Llama3.2-VL-90B-Ins 4.14 0.00 23.62 0.00 7.23
Qwen2.5-VL-72B-Ins 28.51 16.65 45.61 0.00 47.77
# Doc2SAR Framework
GPT-4o 67.68 59.58 79.88 60.58 75.69
Llama3.2-VL-90B-Ins. 64.32 59.57 69.92 55.37 72.18
Qwen2.5-VL-72B-Ins. 75.55 70.67 82.52 64.07 81.03
Qwen2.5-VL-7B-Ins. 73.76 68.27 81.95 59.03 82.62
Ours 80.78 74.09 90.99 65.33 90.08
Overall Performance
The results in Table 1 reveal two key findings regarding overall performance. First, a stark performance gap exists between our synergistic framework and end-to-end methods. End-to-end models struggle significantly, with even powerful models like GPT-4o achieving only 29.30% table recall , and Llama3.2-VL-90B-Ins yielding almost no rational results (4.14%). In contrast, integrating these models into our Doc2SAR framework, which uses domain-specific tools for Layout Detection and OCSR, yields dramatic performance boosts. Second, targeted fine-tuning is critical for state-of-the-art performance. Our SFT-enhanced model (Ours) achieves the highest overall table recall of 80.78%. Notably, our model, fine-tuned on the Qwen2.5-VL-7B-Ins., significantly outperforms the much larger model Qwen2.5-VL-72B-Ins. (80.78% vs. 75.55%).
Performance by Document Type.
Analyzing the results by document type shows our framework’s robustness across different formats. The framework approach is particularly effective for patents, which often feature dense text and complex table structures. Furthermore, the SFT process delivers a pronounced improvement on literature articles, raising the recall from 81.95% (for Qwen2.5-VL-7B-Ins) to 90.99% (Ours). This is likely because the PubTabNet data used for fine-tuning aligns well with the typical table layouts found in scientific literature.
Sub-Task Analysis.
The sub-task breakdown uncovers the fundamental weaknesses of the end-to-end approach. All end-to-end models completely fail on the Coreference Recognition sub-task, scoring 0.00% table recall. This failure stems from the inherent inability of current general-purpose MLLMs to perform accurate OCSR and subsequently link the recognized structures to their correct coreferences. Our synergistic framework, by incorporating a dedicated OCSR tool, effectively solves this bottleneck. A similar performance gap is also observed in the Table Extraction sub-task, further validating the necessity of a modular, specialized approach.
4.3 Key Components Analysis: The impact of SFT
To precisely quantify the benefits of SFT, we conducted a granular analysis on the two key MLLM-driven components: Coreference Recognition and Table Structure Recognition. These experiments isolate the impact of SFT by comparing our fine-tuned model against its base architecture and the zero-shot GPT-4o baseline, demonstrating how targeted training enhances capabilities that are critical to our framework’s overall success. The detailed results are presented in Table 2 and Table 3.
Coreference Recognition.
We evaluate this task on our custom Coreference Val Set using Recall (%) as the metric. Our SFT-enhanced model (Ours) achieves an overall recall of 76.40%, significantly outperforming both GPT-4o (53.40%) and its base model, Qwen2.5-VL-7B-Ins. (36.60%). The improvement is most pronounced in the unstructured Mol-outside-Table scenario (70.45% vs. 38.81% for GPT-4o). In the more structured Mol-inside-Table setting, our model also remains highly competitive (95.76%), demonstrating its robustness across different layouts. This confirms that domain-specific SFT is crucial for handling the varied contexts in which molecules appear.
Table 2: Performance comparison on the Molecular Coreference Recognition task (Recall %).
Model Recall (%)
Overall Mol-inside-Table Mol-outside-Table
GPT-4o 53.40 99.39 38.81
Qwen2.5-VL-7B-Ins. 36.60 88.48 17.91
Ours 76.40 95.76 70.45
Table HTML Extraction.
We evaluate this task on the standard PubTabNet Val Set using the TEDS score (%) [29]. The results again highlight the power of SFT. Our model achieves a superior overall TEDS score of 58.29%, compared to 48.05% for GPT-4o and 43.51% for the base model. Its advantage is largest on tables with Complex Layouts (54.44% vs. 43.31% for GPT-4o), which involve challenging merged cells. Even on Standard Grids, our model maintains a clear lead (61.72% vs. 52.57% for GPT-4o), indicating an all-around improvement. This enhanced structural understanding is a key factor in the high performance of our Doc2SAR framework.
Table 3: Performance on the Table HTML Extraction task, measured by average TEDS score (%).
Model Avg. TEDS score (%)
Overall Standard Grids Complex Layouts
GPT-4o 48.05 52.57 43.31
Qwen2.5-VL-7B-Ins. 43.51 49.24 37.21
Ours 58.29 61.72 54.44
5 Case Study
To provide qualitative evidence of our framework’s real-world efficacy, this section presents two representative case studies from diverse scientific documents. These cases demonstrate Doc2SAR’s ability to accurately process varied layouts and correctly associate molecular structures with their activity data, even under challenging conditions.
Figure 5: Case study in literature 10.1021_acsmedchemlett.3c00392. Note that the red and blue boxes (masks) in the case are only for auxiliary visualization display and are not included in the document. Doc2SAR correctly performs intra-page association, linking a molecular structure to its corresponding activity data table on the same page.
Figure 6: Case study in patent CN105793260B. Note that the red and blue boxes (masks) in the case are only for auxiliary visualization display and are not included in the document. Doc2SAR demonstrates its robustness by performing a long-range, cross-page association, linking a molecule on Page 34 with its activity data in a table on Page 70.
The first case, illustrated in Figure 5, involves a typical scientific literature article where the molecular structure and its corresponding activity data are located on the same page. Our framework successfully performs this intra-page association, correctly identifying the molecule, locating its entry in the activity table, and extracting the complete data row.
The second case, shown in Figure 6, highlights a more challenging scenario from a Chinese patent. The molecular structure is found on Page 34, while its activity data is in a table 36 pages later on Page 70. This large, cross-page span poses a significant challenge for monolithic end-to-end models. In contrast, our modular Doc2SAR framework excels, robustly linking the disparate information via its coreference-based merging strategy. This case also showcases the framework’s versatility in handling non-English documents.
For a more comprehensive analysis, including additional successful extractions and a discussion of challenging failure cases, please refer to Appendix D.
6 Related Work
Chemical Information Extraction.
Early rule-based systems could recognise chemical names or parse fixed reaction templates, but generalised poorly to diverse layouts. Recent efforts have shifted toward neural, modality-specific extractors. For patents, the ChEMU2022 shared task released a benchmark for chemical NER and event extraction with layout cues [30]. Transformer OCSR models—DECIMER.ai with image–to–sequence decoding [31], MolScribe’s image-to-graph generation [32], and MolGrapher’s detection-plus-graph pipeline [33]—now achieve >90% top-1 exact match on curated sets. PatCID further demonstrated large-scale extraction of 80 M structure images from patents by combining segmentation with OCSR [34]. However, these systems operate on a single input type (text or image) and therefore cannot recover cross-modal SARs.
Domain Large Language Models.
Large language models are increasingly adapted for chemistry and materials science. ChemCrow [35] couples GPT-4 with cheminformatics tools to plan multi-step syntheses; Vangala et al. [36] fine-tuned GPT-J to mine patent reactions, boosting recall by 26 % over grammar-based baselines; FuncFetch [37] uses GPT-3.5 to extract >20 k enzyme–substrate pairs with high precision; and SynAsk [38] combines a knowledge base with a GPT-derived model for interactive retrosynthesis and reaction-condition suggestions. Although these LLMs excel at text-centric tasks, evaluations show degraded accuracy once molecular diagrams or numeric tables must be aligned with textual mentions [39]. Doc2SAR addresses this limitation by integrating fine-tuned multimodal LLMs with domain-specific tools to link text, figures and tables in documents.
Multimodal Documents Understanding.
Vision–language document models fuse layout and visual cues: LayoutLMv3 [40] unifies text–image masking for pre-training, UReader [41] adopts an OCR-free encoder for long documents, and DocOwl 1.5 [42] refines multimodal alignment for page-level QA. Benchmarks such as UniMMQA [43] and SPIQA [44] probe reasoning across text, tables and figures, yet contain few chemical diagrams and no ground-truth SAR links. Therefore, DocSAR-200 contributes the first dataset in which molecular structures and activity tables are jointly annotated, and Doc2SAR provides a framework that demonstrably outperforms general VLMs on this document-level SARs extraction.
7 Conclusion
In this work, we addressed the critical challenge of automated SAR extraction from complex scientific documents. We introduced DocSAR-200, a comprehensive new benchmark comprising 200 diverse documents, specifically designed to drive progress in this domain. Concurrently, we proposed Doc2SAR, a novel synergistic framework integrating domain-specific tools with advanced MLLMs which enhanced by SFT. Our experiments demonstrate that the SFT-enhanced framework achieves a state-of-the-art Table Recall of 80.78%, significantly outperforming the strongest end-to-end baseline by a remarkable 51.48% over GPT4o. This result highlights the superiority of our modular design by demonstrating its ability to excel at fine-grained sub-tasks where general models falter, while also affirming the crucial role of SFT in adapting models for precise, domain-specific tasks. By providing robust tools and a standardized framework, our work helps accelerate the transition from manual curation to automated knowledge discovery in the sciences.
While Doc2SAR sets a new standard, we acknowledge current limitations, such as a primary focus on English documents and the need for curated data for SFT, which define clear avenues for future work:
  • Benchmark Expansion: We plan to expand DocSAR-200 with more languages (e.g., Chinese, German), diverse scientific domains, and more challenging annotation types.
  • Methodological Advancements: We will investigate hybrid synergistic models and explore more data-efficient fine-tuning techniques to reduce dependency on supervised data.
  • Broader Applications: We aim to adapt our framework to other critical scientific tasks, such as extracting experimental protocols, to foster a more comprehensive AI-driven ecosystem.
We believe that DocSAR-200 will serve as a valuable resource for the community, spurring the development of more sophisticated and accurate models for scientific document understanding. Doc2SAR provides a strong baseline and a flexible architecture that can be built upon for future research in this rapidly evolving field.
Appendix A Prompts
This section details the specific prompts within our Doc2SAR framework. These prompts were carefully designed to guide the models in performing their specialized sub-tasks: extracting activity data from table images into HTML format, and recognizing molecular coreferences from molecular structure images, distinguishing between cases where the structure is found outside or inside a table.
Prompt for extracting activity data from table image System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to extract activity data from an image in tabular HTML format.
User Message:
First, analyze the table content from the provided image and convert it into an HTML format. The HTML should include all rows and columns, representing the table structure exactly as it appears in the image.
Use the following guidelines:
1. Determine if the table has a header row. If the first row is a header, represent its cells with `<th> `elements. Otherwise, represent all rows (including the first row) with `<td> `elements.
2. Include a <table> element containing all rows (`<tr> `) and cells (`<th> `for headers, `<td> `for data).
3. Retain the exact content of each cell, including special characters, numerical values, text and even empty cell. 4. If a cell contains molecular structure images, replace it with the token "[mol]".
5. If the table contains merged cells, use appropriate attributes like `rowspan `or `colspan `to reflect the merging.
6. Pay attention to the following content: (1) Attribute: [’EC50’, ’IC50’, ’Ki’, ’Kd’, ’pKd’, ’TD50’, ’Ti’, ’TC50’]
(2) Unit: [’uM’, ’nM’, ’%’, ’kcal/mol’] If none of these Attribute keywords are found, it is not activity table. Your entire output should be the exact string: <table>None</table>
Example HTML structure:
```html
<table>
<tr>
<th>xxx</th>
<th>xxx</th>
<th>xxx</th>
</tr>
<tr>
<td>xxx</td>
<td>x</td>
<td>x</td>
</tr>
<tr>
<td>xxx</td>
<td>x</td>
<td>xx</td>
</tr>
</table>
[Image of a table may contains activity data]
Prompt for recognizing coreference from molecular structure image (In Case of Outside the table) System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to recognize the coreference from a molecular structure image in json format.
User Message:
I have a molecular structure image **not inside a table**.
Your task is to **extract only the molecular coreference** associated with the molecule. The coreference is typically a unique identifier (e.g., numeric, alphabetic, or alphanumeric code) found in the image.
Important Instructions:
1. **Focus on coreference priority**:
- **First, check the left side of the molecule** for a valid coreference.
- **If the left side is empty or invalid, check the right side.**
- Do **not** check inside any table.
2. **Ignore the molecular structure itself**:
- Do not output SMILES, InChI, or any chemical structure representation.
- Focus only on extracting the **molecular coreference**.
3. Strictly Follow The JSON Output Format
Example Json structure:
```json
{{"compound_id": "<Molecular coreference or [None]>"}}
[Image of a molecular structure contains corresponding coreference]
Prompt for recognizing coreference from molecular structure image (In Case of Inside the table) System Message: You are an expert in the field of pharmaceutical chemistry, and your task is to recognize the coreference from a molecular structure image in json format.
User Message:
I have a molecular structure image located **inside a table**.
Your task is to **identify only the molecular coreference** in the image, which is typically a unique identifier (e.g., a numeric, alphabetic, or abbreviations) associated with the molecule in the image.
Important Instructions:
1. **Priority of Coreference Extraction**:
- The molecular coreference **is usually inside the same table cell** as the molecular image.
- If the table cell is empty or ambiguous, check nearby table cells.
- Do **not** look outside the table.
2. **Ignore the molecular structure itself**:
- Do not output SMILES, InChI, or any chemical structure representation.
- Focus only on extracting the **molecular coreference**.
3. Strictly Follow The JSON Output Format
Example Json structure:
```json
{{"compound_id": "<Molecular Coreference or [None]>"}}
[Image of a molecular structure contains corresponding coreference]
Appendix B Annotation Details and Quality Control
The DocSAR-200 annotation process was designed to ensure high fidelity and reproducibility. Each document was processed through a two-stage pipeline: first, GPT-4o’s multimodal capabilities [17] were employed to extract preliminary tabular information (e.g., compound names, activity values) into structured formats. In parallel, molecular structure images were located using coreference links or positional heuristics.
Subsequently, trained annotators reviewed and corrected all entries, ensuring accurate alignment between molecules, labels, and extracted tables. Molecular images were parsed into SMILES representations via open-source OCSR tools, such as MolParse [20], followed by human validation.
Challenge considerations. Several challenges were addressed during annotation. Multi-page layouts—common in patents—were handled using persistent compound IDs and bounding box coordination to track dispersed molecular and activity references. Markush structures, which cannot be resolved into single molecules, were excluded. For multilingual documents (English, Chinese, Japanese), language-specific tokenization and layout normalization were applied. Native speakers ensured semantic consistency across languages.
Validation protocols. All SMILES were validated using RDKit [45] to ensure chemical correctness and canonicalization. Heuristic filters removed malformed structures and disconnected fragments. Coordinates and page indices were manually verified to prevent drift or misalignment.
Annotation statistics. The finalized dataset contains 3329 valid SMILES across 200 documents (98 patents, 102 research articles) in three languages. The annotation process required over 160 hours of expert effort.
These measures collectively ensure that DocSAR-200 offers a reliable, high-quality benchmark that reflects the real-world complexity of SAR extraction.
Appendix C SFT Dataset Construction Details
SFT dataset is meticulously constructed from a combination of real-world and synthetic molecular images, designed to enhance the model’s capabilities in recognizing molecular coreferences. To create a unified training corpus, all samples are processed into one of two standardized layouts: an in-table layout, which captures the entire table row containing the target molecule, and an out-of-table layout, where the molecule image is cropped with fixed padding to preserve essential surrounding context.
The real data samples originate from a diverse collection of patents and scientific PDFs, reflecting authentic, "in-the-wild" scenarios. Our semi-automated annotation pipeline begins with UniFinder to detect the bounding boxes of molecules, followed by PyMuPDF [46] to crop the corresponding images based on their layout type. An initial annotation pass is performed by Qwen2.5-VL-7b-Instruct [24] using domain-specific prompts (see Appendix A), with all generated labels subsequently undergoing a rigorous two-person verification process to ensure accuracy. To further augment the dataset, especially for covering edge cases and creating challenging negative samples, we generate synthetic data. These samples are created by rendering molecular structures from PubChem [26] into both layouts. Their labels are randomly generated based on real-data formats, and for out-of-table samples, label positions are programmatically assigned using OpenCV [47] to approximate empirical distributions (e.g., bottom: 50%, top: 30%).
The final consolidated SFT dataset comprises a total of 9,790 samples. This includes 4,594 real-world samples, of which 1,747 are in the in-table layout and 2,847 are out-of-table. The dataset is further enriched by 5,196 synthetic samples, composed of 510 in-table instances and a large number of out-of-table hard cases (4,686), specifically designed to enhance the model’s robustness against complex visual contexts. Figure 7 provides a visual summary of these distributions.
Figure 7: Statistical composition of the SFT dataset for the Mol ID agent. The chart displays absolute sample counts for: (i) In-table and out-of-table components within Real Data, Synthetic Data, and the Overall Dataset. (ii) A direct comparison of total Real Data versus total Synthetic Data under ’Source Totals’.
Data Quality Control.
All real-data annotations underwent double labeling and third-party adjudication, supported by automated checks on field integrity, SMILES validity, and spatial alignment. For synthetic data, layout logic and label formats were systematically validated to ensure consistency with observed real-world patterns. This combination of rigorous human review and controlled synthesis maintained both precision and diversity within the SFT corpus.
Appendix D Case Study
To further illustrate the capabilities and limitations of Doc2SAR, this appendix presents additional case studies. We showcase two representative examples below. The first is a successful extraction from a patent (Figure 8), demonstrating the framework’s robustness in accurately processing well-structured tables with embedded molecular images. In contrast, the second case presents a failure from a scientific article (Figure 9) that highlights a key challenge: the misidentification of molecules within complex reaction schemes. This transparent analysis provides insight into both the system’s current strengths and its areas for future development.
Figure 8: A successful extraction case from patent CN102574816A. Doc2SAR accurately processes the provided table, correctly extracting the molecular structures, their coreference IDs (e.g., A35, A36), and associated activity values into a clean, structured output.
Figure 9: A challenging failure case from literature CN102482273B. The system incorrectly extracts the SMILES string for the molecule with coreference ’1’. This misidentification may be attributed to the complexity of the reaction scheme diagram, where the simple label ’(1)’ refers to the initial reactant but is located within a visually complex multi-step synthesis context, leading to an incorrect association.
Appendix E Web Application
To demonstrate the practical utility and facilitate user interaction with our Doc2SAR system, we have developed an intuitive web application. This platform serves not only as a showcase for the end-to-end extraction capabilities of Doc2SAR but also as a tool for users to verify results and understand the system’s performance on their own documents. The application allows users to easily upload PDF documents (scientific articles or patents) and submit them for automated molecular activity information extraction.
Upon completion of a task, the extracted information is presented in a clear, tabular format, as illustrated in Figure 10(a). Each row typically corresponds to an extracted molecule-activity pair, displaying the molecule coreference, its SMILES representation (potentially with a rendered 2D chemical structure image), and the associated activity values (e.g., IC50, EC50, units). A key feature of the web application is its traceability functionality. Users can click on an extracted SMILES string or an activity data point within the table, and the system will navigate to and highlight the corresponding source region in the original PDF document, as shown in Figure 10(b). This direct linking between the extracted data and its origin within the document is crucial for transparency, allowing for quick verification of the extraction accuracy and building user confidence.
Furthermore, recognizing that automated systems may occasionally require refinement, the web application incorporates an interactive editing interface (Figure 10(c)). Users can directly modify extracted coreferences, SMILES strings, or activity values. For SMILES correction, the interface supports direct text editing, drawing the chemical structure using an embedded molecular editor, or even uploading a corrected molecular image from which the SMILES can be re-derived. Finally, the curated and validated extraction results can be conveniently downloaded by the user in CSV format for downstream analysis and integration into their research workflows (Figure 10(d)). A live demonstration of the web application is accessible at https://uni-finder.dp.tech/molecular-extract.
Figure 10: Screenshots of the Doc2SAR web application: (a) Tabular display of extracted molecular activity data. (b) Example of traceability, highlighting the source of an extracted SMILES string in the original PDF. (c) Interface for interactive editing and correction of extracted results. (d) Functionality for batch uploading files and downloading the extracted results.